### v4

# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

Abstract
Optimal processes in stochastic thermodynamics are a frontier for understanding the control and design of non-equilibrium systems, with broad practical applications in biology, chemistry, and nanoscale/mesoscale systems. Optimal transport theory and thermodynamic geometry have emerged as optimal control methodology, but they are based on slow-driving and close to equilibrium assumptions. An optimal control framework in stochastic thermodynamics for finite time driving is still elusive. Therefore, we solve in this paper an optimal control problem for changing the control parameters of a discrete-state far-from-equilibrium process from an initial to a final value in finite-time. Optimal driving protocols are derived that minimize the total finite-time dissipation cost for the driving process. Our framework reveals that discontinuous endpoint jumps are a generic, model-independent physical mechanism that minimizes the optimal driving entropy production, whose importance is further amplified for far-from-equilibrium systems. The thermodynamic and dynamic physical interpretation and understanding of discontinuous endpoint jumps is formulated. An exact mapping between the finite-time to slow driving optimal control formulation is elucidated, developing the state-of-the-art of optimal transport theory and thermodynamic geometry, which has been the current paradigm for studying optimal processes in stochastic thermodynamics that relies on slow driving assumptions. Our framework opens up a plethora of applications to the thermodynamically efficient control of a far-from-equilibrium system in finite-time, which opens up a way to their efficient design principles.
1 Introduction
1.1 Background and Motivation
The framework of Stochastic Thermodynamics (ST) has developed the thermodynamic understanding of microscopic/mesoscopic systems [1, 2, 3, 4, 5], thermodynamically-consistent coarse-grained macroscopic systems [5, 6], and the trade-off between thermodynamic length, dissipation, and fluctuations [7, 8]. In ST, the entropy production rate (EPR) quantifies the thermodynamic dissipation cost and constrains the dynamics of non-equilibrium systems. For instance, the fluctuation relation (FR) and the thermodynamic kinetic uncertainty relation (TKUR) have revealed fundamental thermodynamic laws, valid beyond the second law of thermodynamics [1, 2, 3, 4, 5, 9, 7, 8]. Recently, a new paradigm of optimal control problems has emerged in ST. Here, the optimal control problem is loosely defined as transforming an initial state (control parameter) into a final state (control parameter) in a finite time, to compute the optimal driving protocols (change of state/control parameters) that minimize the thermodynamic dissipation. Three classes of methodologies have emerged to study optimal control.
First, optimal processes that minimize finite-time dissipation have been investigated for a stochastic particle in a harmonic trap [10]. Surprisingly, the finite-time optimal protocols were found to exhibit discontinuities at the initial and final times, namely the ‘kinks’. Recently, this mechanism has been understood as a mathematical artefact of the imposed boundary condition [11], and was inspired by model and constraint-specific methodologies developed for quantum systems [12], that are valid under certain approximations, and do not hold for generic cases  [^1] These different methodologies are related to different optimal control formulations of classical stochastic systems discussed subsequently, but they are not necessarily the same. Fundamental differences between them can be drastic, despite their case-specific applicability. First, their specific aims differ significantly depending on what they are designed for, mainly the dynamical aspects; see Ref. [12]. Second, the non-equilibrium notion of entropy production rate itself is not necessarily well defined for quantum thermodynamic systems, but it is of paramount importance for classical stochastic thermodynamic systems. As a result, these quantum methodologies strictly focus on the optimization of dynamical aspects, such as constraints on the dynamical evolution of the quantum state or operators, or on their eigenstates.   These methods justify the importance of dynamical aspects over thermodynamic aspects for finite-time optimal processes, arguing that dynamics become more significant than thermodynamics. However, by construction, the most fundamental principles of statistical physics imply that thermodynamics constrains the dynamics, and not the other way around. Therefore, even the dynamics of finite-time optimal processes ‘must’ be constrained by thermodynamics. We rigorously prove this within our framework. . However, these works fail to capture the physical and thermodynamic origin of the ‘kinks[10, 11]. The universality of ‘kinks’ have been observed in some computationally solvable models [14, 15, 16, 17]. Despite a few model-specific studies, analytically computable solutions for other models are lacking, which has created a void in the theoretical physical understanding/framework to attribute this phenomenon.
Second, thermodynamic length defines the distance in the control parameter space of a model and connects it to the thermodynamic dissipation cost [18, 19, 20, 21, 22, 23]. This Riemannian geometric structure in the control parameter space has been exploited to compute optimal driving protocols and formulate the framework of thermodynamic geometry, which is valid in the slow-driving limit. Within this framework, a geodesic is a minimum-distance path between the initial and final control parameters and is equivalently the optimal driving protocol [23, 24]. Thermodynamic geometry has advantages because of its practical applicability. In particular, metric tensors are numerically/experimentally computed for sophisticated models and systems [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 34]. Therefore, because of its practical applicability, thermodynamic geometry has been rigorously studied in comparison to the first methodology. However, thermodynamic geometry has two major drawbacks. First, it relies on the slow-driving approximation, which makes it suboptimal for finite-time optimal driving processes, where the driving time is small and the slow-driving assumption is inherently violated. See Ref.[37] for the experimentally verified violation. Second, it lacks the ‘kinks’ in optimal driving protocols, rendering it inconsistent with exact analytical solutions in Ref. [10, 11, 14, 17, 15, 16]. Thermodynamic geometry is a sub-case of the general mathematical framework known as information geometry; however, even information geometry does not address these drawbacks [38].
Recently, a third optimal control methodology has emerged in ST, which uses a mathematical framework: optimal transport theory [39, 40, 41, 42]. The insights from optimal transport theory have been incorporated into ST [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The mapping between ST and optimal transport theory relies on the equivalence between EPR in ST and the Wasserstein distance in optimal transport theory (an information-theoretical distance measure between probability distributions). The optimal transport map corresponds to optimally changing an initial probability distribution to a final probability distribution; equivalently, the optimal driving protocol is obtained here. Hence, the optimal control in optimal transport theory assumes ‘full’ control of the probability distribution defined in the state-space; in comparison, thermodynamic geometry assumes a parametric control defined in the control parameter space. Due to its inherent formulation as an optimization problem, optimal transport theory has a broad range of applications statistical machine learning and computer science [57], to name a few, computer vision, linguistic, signal, and image representation. It is also used to solve nonlinear differential equations [42], which are of paramount importance in physics [39], biology, economics, and fluid mechanics [40].
Despite its multitude of successful applications, optimal transport theory has three major drawbacks. First, despite the numerical applicability of optimal transport theory, exact analytical solutions are not available except for Gaussian systems, a statistical approximation that does not necessarily hold for finite-size systems in ST prone to non-Gaussian fluctuations [7, 8]. Second, the quadratic dependence of EP on the transition affinity, it is an assumption that holds for systems close-to-equilibrium (cEQ) [7, 8, 5, 6], and leads to a massive underestimation of EP for far-from-equilibrium (fEQ) systems [7, 8]. Third, when the driving time to reach the initial to final state is finite, driving time is a resource to be optimized over, optimal transport theory does not take this constraint into account, which results in its suboptimal performance for a finite-time optimal process, since optimal transport theory is built upon on the slow driving assumption. Therefore, the results obtained using optimal transport theory are inconsistent with the model-specific exact analytical results from Ref.[10, 14], the ‘kinks’ are absent even for the most simple models; see Ref.[58] for a comparison.
The total EP has three linearly independent contributions, namely driving work, excess EP, and housekeeping EP [5, 8]. They physically correspond to dissipation due to total free energy, relaxation towards the Boltzmann distribution, and sustaining nonconservative forces, respectively. However, in the work discussed so far, the focus of optimization has been on driving work or excess EPR. The optimization of the housekeeping EPR has been completely lacking, which quantifies the thermodynamic cost of sustaining non-equilibrium currents. In addition, the focus has been on continuous-state systems, instead of discrete-state systems. Where, due to the non-quadratic dependence of the EPR on driving affinity in discrete-state systems, cEQ optimal control methods are physically less relevant, as they assume a quadratic dependence of the EPR on driving affinity and neglect housekeeping EPR [7, 8]. Despite attempts to understand the finite-time optimal processes consistent in ST, inconsistencies and discrepancies persist, and a coherent unified framework and understanding of the finite-time optimal control in ST is lacking.
Due to their better computational aspects, variational formulations of physical processes have been utilized to study the dynamics of non-equilibrium systems, for instance, phase transitions, first-passage times, and metastability in stochastic systems [59, 60, 61, 62, 63, 64, 65, 66], efficient numerical optimization problems in machine learning [67, 68, 69], and recently explored in ST [70, 71, 72, 73, 74, 75]. Recently, an exact variational formulation for fEQ discrete-state processes has been formulated, namely, the ‘Minimum action principle’ (MinAP) [7, 8], an exact canonical ensemble analog of non-equilibrium discrete-state systems, that unifies FR and the non-quadratic formulation of TKUR within a single framework [7, 8].
1.2 Summary of main results
A Overview
Here, based on MinAP [7, 8], we formulate the generalized finite-time optimal control (GFTOC) framework, valid for any discrete-state fEQ system, by exploiting the Riemannian geometric structure in the control-parameter space. We compute the optimal finite-time driving protocols that minimize the total driving EP, which exhibit the ‘kinks’. GFTOC unifies the optimal control of slowly driven and finite-time processes, as we prove an exact geometric mapping between them. Importantly, since MinAP assigns a thermodynamic cost for any transition-path realization, a thermodynamic cost for sustaining the ‘kinks’ is quantified (a boundary term in driving time), namely the thermodynamic shock. Thereby, the ‘kinks’ reduce the thermodynamic cost of driving (a bulk term in driving time), a ‘geometric shortcut’, since the ‘kinks’ reduce the distance to be travelled in the control-parameter space. The thermodynamic interplay between the boundary and the bulk terms of driving EP gives rise to the formation of the ‘kinks’ as a dynamic consequence, and the restoration of ‘the finite-time geometric speed limit’, an inherent physical timescale associated with the optimal driving process, ‘a time-dissipation-geometry tradeoff relation’. The dual manifestation of the bulk–boundary terms of driving EP has a physical interpretation analogous to work–heat, but formulated here for the finite-time optimal driven process in ST.
We demonstrate the applicability of GFTOC through several examples and address consistency issues arising from existing model-specific optimal control formulations. This leads to a unified and universal theoretical framework for optimal control in ST. Our approach provides a threefold realization of the minimum average entropy production (MinAP) principle in ST via the non-quadratic TKUR, fluctuation relations (FR), and GFTOC, corresponding to thermodynamic inference, partial control, and full control descriptions, respectively [7]. The full-control description enabled by GFTOC allows for the thermodynamically efficient design and control of finite-size stochastic systems in finite driving time, where ST has been experimentally validated. The stochastic-thermodynamic insights of the ‘kinks’ in GFTOC advance optimal transport theory and thermodynamic geometry beyond the slow-driving and close-to-equilibrium assumption. Since information geometry accounts only for excess EP and lacks the possibility of kinks [38], GFTOC further advances the state-of-the-art in information geometry.
B Technical outline
We explicitly define five shortcomings of the existing cEQ optimal control methodologies implemented in ST so far.
  • (S1):
    Violation of the Riemannian geometric structure in control parameter space.
  • (S2):
    Exclusion of the housekeeping EPR.
  • (S3):
    Use of quadratic cEQ dissipation functions for the EPR.
  • (S4):
    Constraint of finite driving time.
  • (S5):
    Lack of thermodynamic and dynamic origin and interpretation of kinks.
The exact model-specific results in Ref.[10] focus on resolving $(\text{S}4)$ but do not address $(\text{S}1),(\text{S}2)$ and $(\text{S}5)$. Despite addressing $(\text{S}2)$, optimal transport theory is prone to $(\text{S}1)$, $(\text{S}3)$, and $(\text{S}4)$ 222optimal transport theory introduces an incremental development $(\text{S}3)$ that addresses $(\text{S2})$, however, it is still a fundamental shortcoming for fEQ systems, see Refs.[7, 8]. Therefore, by construction, methodologies with $(\text{S}2)$ are also prone to $(\text{S}3)$. . thermodynamic geometry aims to address $(\text{S}1)$, but remain susceptible to $(\text{S}2)$ and $(\text{S}4)$  [^3] By construction, methodologies with $(\text{S}4)$ are also prone to $(\text{S}5)$..
The key results obtained with our framework are:
  • Section˜2 deals with shortcomings $(\text{S}1),(\text{S}2)$ and $(\text{S}3)$ under the slow-driving assumption. The class of optimal fEQ driving protocols given by universal geodesics (eq.˜10) is defined as the linear interpolation in the geodesic space of the control parameters (eq.˜8). The corresponding dynamic, thermodynamic, and timescale quantities (eq.˜8, eq.˜9, and eq.˜15) are defined as optimal driving protocol, optimal driving EP and geometric speed limit, respectively, quantifying the minimal driving distance path, dissipation, and time.
  • Section˜3 addresses shortcomings $(\text{S}4)$ and $(\text{S}5)$ via an exact mapping to the optimal slow-driving control problem in section˜2. The resulting finite-time dynamical, thermodynamic, and timescale counterparts are given by eq.˜21, eq.˜25, and eq.˜27. The thermodynamic cost (eq.˜24) admits a boundary-bulk decomposition, analogous to heat and work, which is physically associated with the generation of kinks and slow driving, respectively. Thereby, restoring the finite-time geometric speed limit (eq.˜27): ‘a time-dissipation-geometry tradeoff relation’.
  • Section˜4 generalizes the analytical framework to two practically applicable settings: multiple control parameters and novel numerical algorithms. Section˜5 applies the framework to case-specific problems, highlighting novel results and solving several previously open problems.
1.3 Setup
Thermodynamically consistent discrete-state systems and graphs. —  We model thermodynamically consistent discrete-state systems using a Markov jump process (MJP) or a chemical reaction network (CRN) represented by a graph [78]. $\rho_{i}$ denotes the probability/density of the state, which is a node of the graph. The set of all discrete states is denoted by $\{i\}$. $\gamma^{\rightleftharpoons}$ denotes the set of forward and backward transitions between states $\rho_{i}$ and $\rho_{j}$, with $J_{\gamma}$ and $A_{\gamma}$ denoting the current and affinity for the transition $\gamma^{\rightleftharpoons}$ between states. $\{\gamma^{\rightleftharpoons}\}$ denotes the set of all bidirectional transitions of the graph. The transitions satisfy the Local detailed balance condition (LDB), $A_{\gamma}=\log{(j_{\gamma}/j_{-\gamma})}=F_{\gamma}-\Delta_{\gamma}E+\Delta_{\gamma}S^{state}$, where $j_{\gamma}$ and $j_{-\gamma}$ are the forward and backward transition currents satisfying $J_{\gamma}=j_{\gamma}-j_{-\gamma}$. The transition affinity $A_{\gamma}$ is decomposed into an external non-conservative driving $F_{\gamma}$, a change in the equilibrium energy functional $\Delta_{\gamma}E$, and a change in the state entropy $S_{i}^{state}=-\log{(\rho_{i})}$ [2]. The energy functional $E(\{\lambda_{E}\})$ is fully controlled using the set of control parameters $\{\lambda_{E}\}$.
The symmetric component of the transition currents is called the traffic and is defined as $T_{\gamma}=j_{\gamma}+j_{-\gamma}$. The transition mobility is defined as $D_{\gamma}=\sqrt{j_{\gamma}j_{-\gamma}}$, and it is time-symmetric, quantifying the amplitude of the transition currents. This allows us to define hyperbolic relations between different basis $\{A_{\gamma},D_{\gamma}\}\to\{J_{\gamma},T_{\gamma}\}:$ $J_{\gamma}=2D_{\gamma}\sinh{(A_{\gamma}/2)}$ and $T_{\gamma}=2D_{\gamma}\cosh{(A_{\gamma}/2)}$. Basis $\{J_{\gamma},T_{\gamma}\}$ and $\{A_{\gamma},D_{\gamma}\}$ formulate the thermodynamic inference and full control descriptions, respectively. This nomenclature is attributed to the affinities $\{A_{\gamma}\}$ and the mobilities $\{D_{\gamma}\}$ being the controllable physical parameters, which physically correspond to controlling the current asymmetry and amplitude, respectively [5].
We introduce a shorthand notation for the state-space column vector $\vec{\rho}=(..,\rho_{i},..)^{T}$ and the current-space vector $\vec{J}=(..,J_{\gamma},..)^{T}$. Therefore, the continuity equation for the transport of probabilities/densities is,
$\begin{split}\partial_{t}\vec{\rho}=\mathbb{S}\vec{J}.\end{split}$ (1)
The stoichiometry matrix $\mathbb{S}$ encodes the topology of the transition space $\{\gamma^{\rightleftharpoons}\}$. It contracts the transition currents $\{J_{\gamma}\}$ to the state-space $\{\rho_{i}\}$. The entries $\mathbb{S}_{i\gamma}$ of $\mathbb{S}$ are $1$ or 1-1 if state $\rho_{i}$ is part of a transition $\gamma^{\rightleftharpoons}$, and the sign convention is decided by the direction of transition currents; otherwise, $\mathbb{S}_{i\gamma}=0$.
Thermodynamic dissipation of graphs. —  The mean EPR for the graph satisfies the bilinear form $\langle\dot{\Sigma}\rangle=\sum_{\{\gamma^{\rightleftharpoons}\}}J_{\gamma}A_{\gamma}$, which physically corresponds to the thermodynamic dissipation being equal to the driving force (affinity) multiplied by the current generated by it. Furthermore, the mean EPR $\langle\dot{\Sigma}\rangle$ for the graph is decomposed into three linearly independent orthogonal contributions [8],
$\begin{split}-\dot{\psi}_{E}&=-\dot{\lambda}_{E}\>\partial_{\lambda_{E}}\psi_{E},\\ \langle\dot{\Sigma}_{E}^{ex}\rangle&=-\sum_{\{i\}}d_{t}\rho_{i}\log{\left(\frac{\rho_{i}}{\rho_{i}^{E}}\right)}=-d_{t}D_{E}^{KL},\\ \langle\dot{\Sigma}^{hk}\rangle&=\sum_{\{\gamma^{\rightleftharpoons}\}}T_{\gamma}^{\perp}F_{\gamma}\sinh{\left(\frac{F_{\gamma}}{2}\right)}.\end{split}$ (2)
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively  [^4] Note that, the choice of energy functional $E$ is more flexible and corresponds to the ‘gauge fixing’ and is usually dictated by the experimental/physical constraint. Therefore, it does not necessarily have to be an equilibrium energy functional, for instance, a non-equilibrium analogue of energy is used for a non-equilibrium steady-state $E=ss$, a Lyapunov functional or rate functional that generates the steady-state distribution. This threefold orthogonal decomposition of mean dissipation has been highlighted in Ref.[8], and was generalized to a fourfold orthogonal decomposition of mean dissipation in Ref.[5] to non-reciprocal systems that break the action-reaction symmetry.. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 8]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 6]. Throughout this work, we use vector ($\vec{*}$) and equivalent set $\{*\}$ representations interchangeably, depending on the context.
Minimum action principle. —  The transition probability measure for discrete-state processes is equal to the exponential of an action $\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]$ [7, 8, 6],
$\begin{split}\mathcal{P}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=e^{-\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]},\end{split}$ (3)
where, the action $\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=\int_{0}^{\tau}dt\>\mathcal{L}^{*}[\{A_{\gamma},D_{\gamma}\}]$ is a time-integral of the effective transition Lagrangian, which equivalently quantifies the effective mean EPR,
$\begin{split}\langle\dot{\Sigma}\rangle=\mathcal{L}^{*}[\{A_{\gamma},D_{\gamma}\}]&=\sum_{\{\gamma^{\rightleftharpoons}\}}2D_{\gamma}A_{\gamma}\sinh{\left(\frac{A_{\gamma}}{2}\right)},\end{split}$ (4)
represented here in the ‘full’ control description using the transition mobility $D_{\gamma}$ and transition affinity $A_{\gamma}$. Equations˜3 and 4 are key results of Refs. [7, 8], and formulate a variational principle for discrete-state processes. In particular, the analytical solution for the dynamics of discrete-state processes is obtained by analytically/numerically solving the variational problem for $\mathcal{L}^{*}[\{A_{\gamma},D_{\gamma}\}]$, namely, the ‘Minimum action principle’ (MinAP). We will exploit the MinAP framework to formulate and solve the GFTOC problem.
2 Slow Driving Optimal Control
We consider a GFTOC problem of slow driving of the effective Lagrangian $\mathcal{L}^{*}$ in eq.˜4 from the initial $A_{\alpha}^{i}$ to the final control parameter $A_{\alpha}^{f}$ in a time $\tau$ that minimizes the EPR required for the driving process, which is implemented through external parametric control of $A_{\alpha}$. Here, the control of transition affinity $A_{\alpha}$ is generally used to denote multiple linearly independent control parameters that can implement this. Three linearly independent cases of affinity control correspond to independently contributed EPR, namely $\partial_{t}\psi_{E},\langle\dot{\Sigma}_{E}^{ex}\rangle$, and $\langle\dot{\Sigma}^{hk}\rangle$, quantified by the corresponding effective Lagrangians and relevant controllable affinities are $\{\lambda_{E}\}$, $\{-\log{\left(\rho_{i}/\rho_{i}^{E}\right)}\}$, and $\{F_{\gamma}\}$, respectively, see eq.˜2.
2.1 Geodesic structure
We consider slow driving of the Lagrangian, which physically implies that the system dynamics and statistics quickly adapt to the instantaneous control parameters. Hence, the slow driving formulation is valid under a timescale separation, precisely, only if the driving timescale is slower than the coupling timescale between the environment and the system or the relaxation timescale of the system, which quantifies how quickly the system adapts to the instantaneously imposed control parameters. Under the assumption of slow driving, we expand the Lagrangian to second-order terms in the rate of driving $\dot{A}_{\alpha}$. The driving Lagrangian reads [20, 21, 22]:
$\begin{split}\mathcal{L}_{drv}^{*}\left[A_{\alpha},\dot{A}_{\alpha}\right]=\frac{1}{2}\partial_{{\alpha}}^{2}\mathcal{L}^{*}\left(\dot{A}_{{\alpha}}\right)^{2},\end{split}$ (5)
such that the total driving EP is $\mathcal{S}_{drv}^{qs}=\Sigma_{drv}^{qs}=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$.
Equation˜5 quantifies the driving kinetic energy, where the mass is given by the instantaneous ‘local curvature’ of the Lagrangian ($\partial_{{\alpha}}^{2}\mathcal{L}^{*}$). We adopt the shorthand notation $\partial_{{\alpha}}^{2}\mathcal{L}^{*}=\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}$, which satisfies,
$\partial_{{\alpha}}^{2}\mathcal{L}^{*}=D_{\alpha}\left[2\cosh{\left(\frac{A_{\alpha}}{2}\right)}+\frac{1}{2}A_{\alpha}\sinh{\left(\frac{A_{\alpha}}{2}\right)}\right]=T_{\alpha}+\frac{1}{4}\mathcal{L}^{*}$ (6)
The fEQ driven systems (with a larger $A_{\alpha}$) exhibit a higher mass, attributed to larger fluctuations ($T_{\alpha}$) and dissipation ($\mathcal{L}^{*}$). Physically, this signifies a higher resistance to the driving for fEQ systems, fig.˜1(a) plots eq.˜6. The fEQ driving is critically slowed due to increased resistance, with a singularity in the $A_{\alpha}\to\infty$ limit, which is caused by divergence of the EPR and traffic. Importantly, the local curvature $\partial_{{\alpha}}^{2}\mathcal{L}^{*}$ generalizes the stochastic Fisher information for any $\mathcal{L}^{*}$, thereby going beyond the existing formulation in information geometry [38]. Choosing excess EP as the total dissipation cost shows the equivalence between mass and stochastic Fisher information for each state, which is the predominantly studied dissipation function in information geometry [38], discussed subsequently below in section˜5.3. We have omitted the potential energy boundary term, namely, the instantaneous EPR, which does not depend on driving speed and time, and corresponds to the unavoidable instantaneous minimum EPR cost, in the limit $\tau\to\infty$. This reduces the slow-driving optimization problem to a minimization problem of the driving kinetic energy (driving EPR). Its solution is the geodesic equation for $A_{\alpha}$,
$\begin{split}\ddot{A}_{\alpha}+\frac{\partial_{{\alpha}}^{3}\mathcal{L}^{*}}{2\partial_{{\alpha}}^{2}\mathcal{L}^{*}}\dot{A}_{\alpha}^{2}=0,\end{split}$ (7)
eq.˜7 generalizes the geodesic equation for the optimal control of fEQ systems that minimizes driving EPR [24]. It equivalently implies $\dot{A_{\alpha}}\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}={v}_{qs}$, where $v_{qs}$ is the quasi-static driving speed along the geodesic in the driving time $\tau$. The minimum action solution for optimal driving is called the geodesic and is denoted by $\mathcal{G}(A_{\alpha})$. It encodes the minimum distance path in the control parameter space, and its analytical form plays a crucial role. By definition, the geodesic implies $\mathcal{G}(A_{\alpha})=\arg\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, and the corresponding minimum EP is $\Sigma_{qs}^{*}=\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, under the initial and final boundary conditions $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$. Thus, $\dot{A_{\alpha}}\propto 1/\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}$ implies that $A_{\alpha}$ has to be driven slower the further the system is from equilibrium. Physically, this means that for the validity of the optimal driving, each infinitesimally small driving timestep $\Delta t$ contributes equally to $\Sigma_{drv}^{qs}$.
Computing the geodesic gives the mapping between driving speed, time, and control parameter, $A_{\alpha}(t)=\mathcal{G}^{-1}(v_{qs}t)$. Here, $\mathcal{G}(A_{\alpha}):A_{\alpha}\to t$ is a function mapping the instantaneous control parameter to time. Its inverse function is $\mathcal{G}^{-1}(v_{qs}t):t\to A_{\alpha}$. Thus, in geodesic space, the driving protocol satisfies the linear solution $\mathcal{G}(A_{\alpha})=c_{0}+v_{qs}t$, with $c_{0}$ and $v_{qs}$ being unknowns to be determined. To this end, slow driving inherently imposes the boundary conditions at the initial and final times. Hence, $\mathcal{G}(A_{\alpha}^{i})=c_{0}$ and $\mathcal{G}(A_{\alpha}^{f})=v_{qs}\tau+c_{0}$, which imply $c_{0}=\mathcal{G}(A_{\alpha}^{i})$ and $v_{qs}=[\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})]/\tau$. This reduces the optimal driving protocol to a linear interpolation in geodesic space between the initial and final control parameters:
$\begin{split}\mathcal{G}(A_{\alpha})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(A_{\alpha}^{i})+\frac{t}{\tau}\mathcal{G}(A_{\alpha}^{f}).\end{split}$ (8)
Hence, ${v}_{qs}=\partial_{t}\mathcal{G}(A_{\alpha})=\frac{1}{\tau}\left[\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right]$. It simplifies $\Sigma_{qs}^{*}=\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)=\inf_{A_{\alpha}}\int_{0}^{\tau}\mathcal{L}_{drv}^{*}dt=\frac{1}{2}\tau({v}_{qs})^{2}=\frac{1}{2}\tau\left(\partial_{t}\mathcal{G}(A_{\alpha})\right)^{2}$. Thus, the minimum slow-driving EP cost reads,
$\begin{split}\Sigma_{qs}^{*}=\frac{1}{2\tau}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2}.\end{split}$ (9)
Figure 1: (a) The curvature of the Lagrangian $\partial_{\alpha}^{2}\mathcal{L}^{*}$ (mass given by eq.˜6) increases exponentially for the higher value of driving affinity $A_{\alpha}$. (b) $\mathcal{G}(A_{\alpha}):A_{\alpha}\to t$. Comparison between $\mathcal{G}^{cEQ}(A_{\alpha})$, $\mathcal{G}^{fEQ}(A_{\alpha})$ and $\mathcal{G}^{lin}(A_{\alpha})$ given by eq.˜10. For the fixed quasi-static driving speed $v_{qs}$, $A_{\alpha}^{f}-A_{\alpha}^{i}=1$ is considered for the cEQ $A_{\alpha}^{i}=1$ and fEQ $A_{\alpha}^{i}=3.5$ and the corresponding $\Delta\tau_{c}$ and $\Delta\tau_{f}$ are plotted; it is a combined pictorial visualization of eqs.˜10 and 15. Due to eq.˜15, the geometric speed limit for the quasi-static slow-driving increases for fEQ systems, an effect attributed to the higher mass required for the driving of fEQ systems.
This simplification allows us to visualize the physical meaning of geodesic space using eqs.˜8 and 9. In particular, the optimal control problem for $A_{\alpha}$ with a varying instantaneous local mass is converted into one with a unit mass using the geodesic map $\mathcal{G}(A_{\alpha})$. Similarly, time is scaled in units of the driving time $\tau$. This mapping converts the non-quadratic optimal control problem in $A_{\alpha}$ space to a quadratic optimal control problem in $\mathcal{G}(A_{\alpha})$ with a scaled unit mass and a scaled unit time. This is a fundamental non-equilibrium scale-invariance relation for the optimal quasi-static driving of fEQ systems, and it delineates the underlying fundamental universality of all quasi-statically driven systems.
2.2 Exact analytical expressions for geodesics
In this section, we compute the exact analytical expressions for the geodesic for the control of transition affinity $A_{\alpha}$ for three different cases. We fix the amplitude of mass to 1 in the $A_{\alpha}\to 0$ limit. The first case is the linear geodesic $\mathcal{G}^{lin}$, which is computed using the equilibrium approximation of mass $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1$. The second case is the close-to-equilibrium geodesic $\mathcal{G}^{cEQ}$, which is computed using the close-to-equilibrium approximation. Up to $O(A_{\alpha}^{2})$ terms of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$, this implies $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1+A_{\alpha}^{2}/4$. The third case is the fEQ geodesic $\mathcal{G}^{fEQ}$, which is computed using the fEQ approximation of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$ using $\cosh{(a)}\approx\sinh{(a)}\approx\frac{1}{2}e^{a}$, implying $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=\frac{1}{4}(A_{\alpha}+4)e^{\frac{1}{2}A_{\alpha}}$.
The geodesic is obtained by solving the ODE $\dot{A}_{\alpha}\sqrt{\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}}=v_{qs}$. In addition, we have imposed the constraint $\mathcal{G}(0)=0$, which corresponds to fixing the integration constant. This allows us to compare different geodesics from the same reference point. The exact closed-form analytical expressions for the linear geodesics $\mathcal{G}^{lin}$, close-to-equilibrium geodesics $\mathcal{G}^{cEQ}(A_{\alpha})$ and far-from-equilibrium geodesics $\mathcal{G}^{fEQ}(A_{\alpha})$ are:
$\begin{split}\mathcal{G}^{lin}(A_{\alpha})&=A_{\alpha},\\ \mathcal{G}^{cEQ}(A_{\alpha})&=\frac{1}{4}A_{\alpha}\sqrt{4+A_{\alpha}^{2}}+\sinh^{-1}\left(\frac{A_{\alpha}}{2}\right),\\ \mathcal{G}^{fEQ}(A_{\alpha})&=2\sqrt{\left(4+A_{\alpha}\right)e^{\frac{1}{2}{A_{\alpha}}}}-\frac{2\sqrt{\pi}}{e}ErfI\left[\frac{1}{2}\sqrt{{4+A_{\alpha}}}\right]\\ &\hskip 14.22636pt-4+\frac{2\sqrt{\pi}}{e}ErfI\left[1\right],\end{split}$ (10)
respectively. Here, $ErfI[x]$ is the imaginary error function. The analytical form of $\mathcal{G}(A_{\alpha})$ plays a key role. $\mathcal{G}^{fEQ}$, $\mathcal{G}^{cEQ}$, and $\mathcal{G}^{lin}$ are plotted in fig.˜1(b). The geodesics satisfy the hierarchy of inequalities,
$\begin{split}\mathcal{G}^{fEQ}(A_{\alpha})\geq\mathcal{G}^{cEQ}(A_{\alpha})\geq\mathcal{G}^{lin}(A_{\alpha}).\end{split}$ (11)
In $A_{\alpha}\to 0$ limit, $\mathcal{G}^{lin}(A_{\alpha}),\mathcal{G}^{cEQ}(A_{\alpha})$, and $\mathcal{G}^{fEQ}(A_{\alpha})$ converge. However, the further the system is from equilibrium, the greater the quantitative difference is observed. The physical implication is a tighter bound on the exact optimal driving EP in eq.˜9, and the deviation of the optimal driving protocol eq.˜8 from a trivial linear interpolation between the initial and final control parameters. This highlights the importance of the exact fEQ geodesics.
2.3 Far-from-equilibrium generalizations of
Optimal transport theory
The geodesic hierarchy eq.˜11 implies the hierarchy $\left(\mathcal{G}^{fEQ}(A_{\alpha}^{f})-\mathcal{G}^{fEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(\mathcal{G}^{cEQ}(A_{\alpha}^{f})-\mathcal{G}^{cEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(A_{\alpha}^{f}-A_{\alpha}^{i}\right)^{2}$, see fig.˜1(b). For a fixed $\tau$, a hierarchy on the driving EP obtained using eq.˜9 reads,
$\begin{split}\Sigma_{qs}^{*fEQ}\geq\Sigma_{qs}^{*cEQ}\geq\Sigma_{qs}^{*lin}.\end{split}$ (12)
$\mathcal{G}^{lin}(A_{\alpha})$ is the special case that has been utilized in optimal transport theory [39, 40], and applied to ST to obtain $\Sigma_{qs}^{*lin}$ as the driving cost of EP, which is quadratic function of the change in the driving affinity [43, 53]. In this aspect, optimal transport theory is a cEQ ‘hybrid’ case that deals with a ‘state-space’ representation, for which the mass for the excess EP is utilized (see section˜5.3)  [^5] There is a minor technical difference that is not of fundamental importance, since the mass defined for a control parameter is the inverse of the mass defined for its dual conjugate; see section 5.3., but a quadratic dependence on the driving-affinity incorporates the housekeeping EPR (see section˜5.4).
However, the computation of $\mathcal{G}$ takes into account the non-quadratic nature of the dissipation function $\mathcal{L}^{*}$, which results in the tightest and exact bound using eq.˜12. Moreover, the right-hand side of eq.˜9 is equal to the square of the $\mathcal{W}_{2}$ Wasserstein distance defined as,
$\begin{split}\mathcal{W}_{2}=\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i}),\end{split}$ (13)
the distance in the geodesic space of the control parameters. This is followed by the definition of $\mathcal{W}_{2}$ and recalling the definition of the geodesic, which physically corresponds to the minimization of distance, and is mathematically equal to the $L^{2}-{norm}$. The definition eq.˜13 of $\mathcal{W}_{2}$ itself is novel, and accounts for the geometric curvature for any control parameter. The existing cEQ optimal transport theory methodology focuses on the control of the state-space. Thus, eq.˜9 leads to equality between the driving EP in ST and the Wasserstein distance in optimal transport theory,
$\Sigma_{qs}^{*}=\frac{\mathcal{W}_{2}^{2}}{2\tau}.$ (14)
Therefore, our optimal control formulation goes beyond the existing quadratic counterparts formulated using the optimal transport theory [43, 53, 54, 58]. Importantly, our slow-driving formulation also reveals the connection between ‘thermodynamic geometry’ and ‘optimal transport theory’ beyond known quadratic formulations. This extends the applicability of ‘optimal transport theory’ to any non-Euclidean (Riemannian) control parameter manifolds.
2.4 Geometric speed limit for slow driving:
time-dissipation-geometry tradeoff relation
The validity of slow driving is a key assumption in the optimal control framework developed so far. Equation˜9 reveals an inherent timescale associated with the driving process. Inverting eq.˜9, we define the quasi-static driving timescale:
$\begin{split}\tau_{qs}^{*}=\frac{\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2}}{2\Sigma_{qs}^{*}}.\end{split}$ (15)
$\tau_{qs}^{*}$ is a quantitative measure of the time required for the given initial and final control parameters ($A_{\alpha}^{i}$ and $A_{\alpha}^{f}$) and the driving EP $\Sigma_{qs}^{*}$. It quantifies the timescale for the violation of the slow-driving assumption. Physically, it is equal to the square of the distance travelled along the geodesic in the control parameter space divided by the quasi-static EP supplied for the driving process.
The quasi-static driving time reveals the fundamental tradeoff between the driving time, dissipation (EP) and geometry for the optimally driven process, and defines the geometric speed limit for the slow-driving process. Therefore, eq.˜15 is a fundamental ‘time-dissipation-geometry tradeoff relation’. Importantly, the slow-driving assumption is violated for any $\tau$ and does not necessarily require $\tau\leq\tau_{qs}^{*}$. For the fixed value $\Delta A_{\alpha}=A_{\alpha}^{f}-A_{\alpha}^{i}$ and a fixed available budget for driving EP $\Sigma_{qs}^{*}$, eq.˜11 implies,
$\tau_{qs}^{*fEQ}>\tau_{qs}^{*cEQ}>\tau_{qs}^{*lin},$ (16)
see fig.˜1(b). Hence, fEQ systems require a larger quasi-static driving time, attributed to critical slowing due to higher mass. This implies that fEQ systems are more prone to breaking the assumption of slow-driving, with prominent consequences discussed subsequently in section˜3.
3 Finite-Time Optimal Control
The finite-time optimal protocol exhibits discontinuous jumps at the endpoints of the protocol, namely, kinks [10, 43, 46, 11]. In contrast, the slow-driving approach misses such jumps [24]. The optimal slow-driving control framework developed in section˜2 relies on the key assumption of a timescale separation between the driving time $\tau$ and the largest inherent relaxation timescale of the system. This implies that the system relaxes instantaneously to the control parameter dictated by the environment, justifying the quasistatic slow-driving assumption. However, finite-time optimal processes could operate on shorter timescales such that $\tau<\tau_{qs}^{*}$, violating the slow-driving assumption. Moreover, the slow-driving assumption is also violated for $\tau\geq\tau_{qs}^{*}$, except in the $\tau\to\infty$ limit. To this purpose, in this section we develop the finite-time optimal control framework, relying on the MinAP combined with the slow-driving optimal control and the geodesic structure.
3.1 Finite-time geodesic structure
In a finite-time driving process, the system realizes that it cannot quasi-statically follow the geodesic. The GFTOC formulation therefore implies that the constraint of the boundary condition on the geodesic has to be relaxed, and needs to be treated as an optimization parameter. We consider $\mathcal{G}(A_{\alpha}^{i*})$ and $\mathcal{G}(A_{\alpha}^{f*})$ as the optimal control parameter values in the geodesic space at $t=0^{+}>0$ and $t=\tau^{-}<\tau$, respectively. Here, $0^{+}$ and $\tau^{-}$ are infinitesimal times after the initial and before the final time, respectively. Hence, the protocol jumps in the geodesic space at the initial and final times are $\Delta_{0+}\mathcal{G}_{\tau}(A_{\alpha})=\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{i})$ and $\Delta_{\tau-}\mathcal{G}_{\tau}(A_{\alpha})=\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{f*})$, where we have chosen the convention $\mathcal{G}(A_{\alpha}^{i})\leq\mathcal{G}(A_{\alpha}^{i*})<\mathcal{G}(A_{\alpha}^{f*})\leq\mathcal{G}(A_{\alpha}^{f})$, which need not be imposed as the optimal solution ensures this hierarchical inequality. The slow-driving is followed along the geodesic from $\mathcal{G}(A_{\alpha}^{i*})$ to $\mathcal{G}(A_{\alpha}^{f*})$ from time $0^{+}$ to $\tau^{-}$. Within the MinAP [7, 8], the thermodynamic EP cost associated with kinks is the boundary term $\Sigma_{bnd}=\int\mathcal{L}_{bnd}^{*}[A_{\alpha}]=\frac{1}{2}\left(\Delta_{\tau-}\mathcal{G}_{\tau}(A_{\alpha})\right)^{2}+\frac{1}{2}\left(\Delta_{0+}\mathcal{G}_{\tau}(A_{\alpha})\right)^{2}$, which penalizes the formation of kinks. The bulk EP to drive along the geodesic from $\mathcal{G}(A_{\alpha}^{i*})$ to $\mathcal{G}(A_{\alpha}^{f*})$ is modified to $\Sigma_{bulk}=\int_{0^{+}}^{\tau^{-}}\mathcal{L}_{bulk}^{*}dt=\frac{1}{2\tau}\left(\mathcal{G}(A_{\alpha}^{f*})-\mathcal{G}(A_{\alpha}^{i*})\right)^{2}$. Since $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$, it follows that $\Sigma_{bulk}<\Sigma_{qs}^{*}$. Hence, the increase in $\Sigma_{bnd}$ due to kinks is compensated by a decrease in $\Sigma_{bulk}$, and their interplay plays a key role in optimal finite-time protocols.
The total EP associated with the GFTOC problem is $\Sigma_{\tau}=\Sigma_{bnd}+\Sigma_{bulk}$. Within the MinAP [7, 8], the minimization of $\Sigma_{\tau}$ requires solving the variational optimal control problem with unknown free parameters $A_{\alpha}^{i*},A_{\alpha}^{f*}$,
$\begin{split}\Sigma_{\tau}^{*}=\inf_{\{A_{\alpha}^{i*},A_{\alpha}^{f*}\}}\left(\Sigma_{\tau}\right).\end{split}$ (17)
The variation with respect to $A_{\alpha}^{i*}$ and $A_{\alpha}^{f*}$ leads to the following set of linear Euler-Lagrange equations $\frac{\delta\Sigma}{\delta\mathcal{G}(A_{\alpha}^{i*})}=0,\frac{\delta\Sigma}{\delta\mathcal{G}(A_{\alpha}^{f*})}=0$, whose reorganization leads to,
$\begin{split}\left(1+\frac{1}{\tau}\right)\mathcal{G}(A_{\alpha}^{i*})-\frac{1}{\tau}\mathcal{G}(A_{\alpha}^{f*})&=\mathcal{G}(A_{\alpha}^{i}),\\ -\frac{1}{\tau}\mathcal{G}(A_{\alpha}^{i*})+\left(1+\frac{1}{\tau}\right)\mathcal{G}(A_{\alpha}^{f*})&=\mathcal{G}(A_{\alpha}^{f}).\end{split}$ (18)
Inverting eq.˜18, its solution reads:
$\begin{split}\mathcal{G}(A_{\alpha}^{i*})&=\frac{1+\tau}{2+\tau}\mathcal{G}(A_{\alpha}^{i})+\frac{1}{2+\tau}\mathcal{G}(A_{\alpha}^{f}),\\ \mathcal{G}(A_{\alpha}^{f*})&=\frac{1}{2+\tau}\mathcal{G}(A_{\alpha}^{i})+\frac{1+\tau}{2+\tau}\mathcal{G}(A_{\alpha}^{f}).\end{split}$ (19)
The corresponding finite-time optimal protocol reads:
$\begin{split}\mathcal{G}_{\tau}(A_{\alpha})&=\mathcal{G}(A_{\alpha}^{i}),\hskip 96.73918ptt=0,\\ &=\left(1-\frac{t}{\tau}\right)\mathcal{G}(A_{\alpha}^{i*})+\frac{t}{\tau}\mathcal{G}(A_{\alpha}^{f*}),\hskip 14.22636ptt\in(0,\tau),\\ &=\mathcal{G}(A_{\alpha}^{f}),\hskip 96.73918ptt=\tau.\end{split}$ (20)
Equivalently, using eq.˜19, the finite-time optimal protocol for $t\in(0,\tau)$ can be expressed using the known control parameters $\mathcal{G}(A_{\alpha}^{i})$ and $\mathcal{G}(A_{\alpha}^{f})$ as
$\begin{split}\mathcal{G}_{\tau}(A_{\alpha})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}(A_{\alpha}^{i})+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}(A_{\alpha}^{f}).\end{split}$ (21)
Equation˜21 is the optimal finite-time transport map. It is equivalent to substituting $t/\tau\to t/(\tau+2)$, $1\to{(1+\tau)}/{(2+\tau)}$, and $0\to{1}/{(2+\tau)}$ in eq.˜8. Physically, optimal finite-time driving is equivalent to total driving time $\tau+2$ with initial and final times $t=1$ and $t=\tau+1$, respectively, leading to jumps in the optimal protocol at the endpoints. This reveals the structural geometric similarities between the finite-time optimal driving and the slow-driving process. Using eq.˜21, the amplitude of kinks is,
$\begin{split}\Delta_{0+}\mathcal{G}_{\tau}(A_{\alpha})=\Delta_{\tau-}\mathcal{G}_{\tau}(A_{\alpha})=\frac{1}{2+\tau}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right).\end{split}$ (22)
Hence, kinks are of equal amplitude in geodesic space, which physically corresponds to equal distribution of the thermodynamic cost associated with kinks. Upon converting back to the control parameter space $A_{\alpha}$ using $\mathcal{G}^{-1}$, kinks have different amplitudes. The higher-mass endpoint has a smaller jump size compared to the lower-mass endpoint. The amplitude of kinks scales as $1/(2+\tau)$, vanishing in the quasistatic limit $\tau\to\infty$. For the fast-driving limit $\tau\to 0$, the optimal protocol follows midpoint interpolation in geodesic space,
$\begin{split}\mathcal{G}_{\tau\to 0^{+}}(A_{\alpha})=\frac{1}{2}\left[\mathcal{G}(A_{\alpha}^{i})+\mathcal{G}(A_{\alpha}^{f})\right].\end{split}$ (23)
Inverting back to the control parameter space gives $A_{\alpha}=\mathcal{G}^{-1}\left(\frac{1}{2}\mathcal{G}(A_{\alpha}^{i})+\frac{1}{2}\mathcal{G}(A_{\alpha}^{f})\right)$ in the limit $\tau\to 0$. $\mathcal{G}_{\tau}(A_{\alpha})$ is plotted in fig.˜2(a) for different $\tau$. Due to $\mathcal{G}(A_{\alpha})>A_{\alpha}$, endpoint jumps are amplified for fEQ systems.
Using the finite-time optimal transport map eq.˜21, the optimal bulk and boundary EP are
$\begin{split}\Sigma_{bulk}^{*}&=\frac{\tau}{2(2+\tau)^{2}}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2},\\ \Sigma_{bnd}^{*}&=\frac{2}{2(2+\tau)^{2}}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2}.\end{split}$ (24)
Equation˜24 reveals the fundamental trade-off between bulk and boundary EP in optimal finite-time processes. In particular, $\Sigma_{bulk}^{*}\propto\tau$ and $\Sigma_{bnd}^{*}\propto 1$, implying that finite-time optimal protocols with kinks are a physical manifestation of the optimization interplay between them. The large-$\tau$ and small-$\tau$ regimes correspond to the domination of $\Sigma_{bulk}^{*}$ and $\Sigma_{bnd}^{*}$, respectively, for the thermodynamic optimization. The total optimal driving EP is, $\Sigma_{\tau}^{*}=\Sigma_{bulk}^{*}+\Sigma_{bnd}^{*}$,
$\begin{split}\Sigma_{\tau}^{*}=\frac{1}{2(2+\tau)}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2}.\end{split}$ (25)
Compared to eq.˜9, $\Sigma_{\tau}^{*}$ is finite in the $\tau\to 0$ limit and smaller than $\Sigma_{qs}^{*}$. They satisfy the non-equilibrium scaling relations $\Sigma_{\tau}^{*}\propto 1/(2+\tau)$ and $\Sigma_{qs}^{*}\propto 1/\tau$, respectively. Importantly, eq.˜25 imposes an upper bound on the driving dissipation,
$\begin{split}\Sigma_{\tau}^{*}\leq\frac{1}{4}\left(\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})\right)^{2},\end{split}$ (26)
which is saturated in the $\tau\to 0$ limit. An analogy of an upper bound eq.˜26 is missing for the slow-driving processes.
Thermodynamically, finite-time optimal driving is equivalent to slow driving with driving time $\tau+2$ instead of $\tau$. This interpretation is consistent with the dynamic mapping between slow-driving and finite-time optimal protocols elucidated using eqs.˜8 and 21. The short-time underestimation from $\Sigma_{\tau}\propto 1/\tau$ scaling has been experimentally observed [37]. In conclusion, we analytically solve the GFTOC problem by revealing its exact geometric connection to its slow-driving counterpart.
3.2 The physical interpretation of kinks and
the role of boundary conditions
We elaborate on the physical interpretation of kinks within our GFTOC framework. In the small driving-time limit, the fundamental principle of ST—the timescale separation between the environment and system degrees of freedom—is violated. This principle normally ensures that the system instantaneously relaxes to the control parameters imposed by the environment. Simultaneously, the boundary conditions enforce a constraint of reaching the final control parameters within the given finite time $\tau$ from the specified initial control parameters. Hence, for small $\tau$, due to violation of the speed-limit bound eq.˜15, the environment cannot instantaneously impose the required control parameters on the system.
Kinks solve this issue, as they reduce the geodesic distance for the finite-time driving process, mathematically expressed as $|\mathcal{G}(A_{\alpha}^{i*})-\mathcal{G}(A_{\alpha}^{f*})|<|\mathcal{G}(A_{\alpha}^{i})-\mathcal{G}(A_{\alpha}^{f})|$. This trade-off between the rush to reach the final point along the geodesic and the inability to do so within time $\tau$ produces kinks. The thermodynamic cost of each kink is $\Sigma_{bnd}^{*}/2$. The system undergoes a ‘thermodynamic shock’ at the initial and final points, generating a spontaneous discontinuous change in the control parameters, which induces a global transition of the system’s state. Within our framework, $\Sigma_{bnd}^{*}$ and $\Sigma_{bulk}^{*}$ can be interpreted analogously to heat and work, respectively: heat corresponds to the instantaneous discrete entropy production cost of a jump, and work to the continuous driving of control parameters. To our knowledge, this provides a novel thermodynamic understanding of the heat–work dual manifestation of the driving EP in finite-time optimal driven far-from equilibrium processes.
3.3 Finite-time geometric speed limit
Figure 2: (a) The finite-time optimal protocol $\mathcal{G}_{\tau}(F_{\alpha})\to t$ is plotted for the different values of $\tau=\{4,1,0.1\}$ with the same initial and final value condition (shown by the dotted blue lines) with $F_{\alpha}^{i}=2$ and $F_{\alpha}^{f}=5$. Its slow-driving counterpart is denoted by ‘qs’. As formulated later in section˜5.5, the vanishing excess affinity is considered because the optimal control of the housekeeping EPR is being solved. Where, eqs.˜8, 21 and 10 are used for the plot and the scaled (with $\tau$) driving time ($t/\tau$) is used. Kinks are of equal amplitude in the geodesic space. (b) The corresponding finite-time non-conservative affinity is obtained using the fEQ geodesic $\mathcal{G}^{fEQ}(F_{\alpha})$ in eq.˜10, which gives the mapping $F_{\alpha}\to t$ in scaled (with $\tau$) driving time ($t/\tau$). Kinks are of unequal amplitude in the affinity space, such that a higher mass point exhibits a smaller amplitude of kinks. (c) The corresponding mapping in $\mathcal{G}_{\tau}(F_{\alpha})\to F_{\alpha}$ space, the initial and final points are denoted by $\star$ and $\blacklozenge$, respectively. Due to the finite driving time constraint, the finite-time optimal protocol eq.˜21 traverses a part of the slow-driving geodesic eq.˜8, a ‘geometric shortcut’, such that the finite-time geometric speed limit eq.˜27 is restored, analogous to the slow-driving geometric speed limit eq.˜15. The choice of discontinuous jumps in the finite-time driving protocol is also constrained by equally distributing the thermodynamic cost ($\Sigma_{bnd}^{*}$ in eq.˜24) between the initial and final endpoints, which generates a ‘Thermodynamic shock’ at the initial and final time.
Using $\Sigma_{bulk}^{*}$ from eq.˜24 and eqs.˜15 and 19, one finds
$\begin{split}\tau_{\tau}^{*}=\frac{\left(\mathcal{G}(A_{\alpha}^{f*})-\mathcal{G}(A_{\alpha}^{i*})\right)^{2}}{2\Sigma_{bulk}^{*}}=\tau.\end{split}$ (27)
This shows that the selection of endpoint jump values $\mathcal{G}(A_{\alpha}^{i*})$ and $\mathcal{G}(A_{\alpha}^{f*})$ is constrained to restore the speed-limit bound eq.˜15 for the finite-time optimal driving process. Physically, this implies that the allowed endpoint jumps are minimal, just sufficient to restore the finite-time speed-limit constraint for any given $\tau$. This highlights a fundamental symmetry: kinks are a non-equilibrium mechanism that circumvents the dynamic constraint of limited driving time quantified by the speed-limit, representing the dynamic counterpart of a ‘thermodynamic shock’.
The upper bound on the total amplitude of kinks, $|\Delta_{0^{+}}\mathcal{G}_{\tau}(A_{\alpha})|+|\Delta_{\tau^{-}}\mathcal{G}_{\tau}(A_{\alpha})|$, a boundary property, can be obtained using the thermodynamic cost of driving and the distance covered in control parameter space (a bulk property). The equal thermodynamic cost $\Sigma_{bnd}^{*}$ for each kink results from minimizing $\Sigma_{bnd}^{*}$ for a constrained total amplitude of kinks. Hence, kinks arise from finite driving-time constraints that may prevent reaching the final state along the geodesic. The finite-time optimal protocol restores eq.˜15 while minimizing the thermodynamic cost of both kinks. Importantly, the slow-driving assumption is broken for any finite $\tau$ and does not require $\tau<\tau_{qs}^{*}$. A continuous change of $\tau$ therefore induces a smooth transition from eqs.˜8 and 9 to eqs.˜20 and 25, with $\tau$ as the relevant control parameter. In conclusion, this mechanism is to be interpreted as a ‘geometric shortcut’. Figure˜2 pictorially summarizes the underlying geometric structure of GFTOC.
3.4 Key features of geodesic
Time-reversal symmetry. —  Equations˜8 and 21 satisfy time-reversal symmetry, as they are invariant under the transformation $\mathcal{G}(A_{\alpha}^{i})\to\mathcal{G}(A_{\alpha}^{f})$, $\mathcal{G}(A_{\alpha}^{f})\to\mathcal{G}(A_{\alpha}^{i})$, and $t\to\tau-t$. This symmetry arises fundamentally from the geodesic structure of the slow-driving process. The symmetric amplitude of kinks extends this symmetry to the finite-time geodesic eq.˜20.
Experimental inference and non-equilibrium scaling relation. —  GFTOC framework implies that the minimum driving EP is obtained by computing the geodesic. If experimental setups can measure the optimal driving EP (thermodynamically constrained to support a minimum dissipation for the driving process), one can pose the inverse problem of reconstructing the geodesic for given initial and final control parameters. The non-equilibrium scaling relation eq.˜25 can be used to infer the analytical shape of $\mathcal{G}(A_{\alpha})$. For the observed set of data $\{(\Sigma_{\tau}^{*},\tau)\}$ at fixed $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$, the collapse of the data yields $\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i})=\sqrt{2(2+\tau)\Sigma_{\tau}^{*}}$. Repeating this procedure for different $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$ reconstructs $\mathcal{G}(A_{\alpha})$ experimentally or numerically. The geodesic shape informs about the non-equilibrium character of the process and the contribution of EP, since it has different analytical dependence on the cEQ and fEQ regimes, as discussed before in section˜2.2 and later in section˜5.
3.5 Comparison and differences to model-specific known results
Finite-time optimal protocols with kinks were analytically computed for a particle in a harmonic trap [10] and recently unified using optimal transport theory [11]. These model-specific results differ from the generalized finite-time framework in three aspects: (i) they optimize the free energy $\psi_{E}$ using control parameters of the equilibrium energy functional $E$, a close-to-equilibrium methodology (ii) the boundary conditions used to obtain kinks differ, and (iii) the thermodynamic cost of kinks is undefined in Ref.[10, 11], leading to the physical interpretation of kinks as a boundary artefact.
We briefly summarize the procedure outlined in [11]. The boundary conditions used in [10, 11] are equivalent to fixing $A_{\alpha}^{i*}=A_{i}$ and computing $A_{\alpha}^{*int}\in(A_{\alpha}^{i},A_{\alpha}^{f})$ that minimize the bulk driving EP under the final value constraint. Thus, $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ is termed as the feasible driving protocol for the given finite-time $\tau$. Then, $A_{\alpha}^{*int}$ is used to compute $A_{\alpha}^{*i}$ and $A_{\alpha}^{*f}$ by introducing counter-diabatic driving terms that shifts the initially computed driving protocol $A_{\alpha}^{i}\to A_{\alpha}^{*int}$ to $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ in driving time $\tau$, and it generates kinks of amplitude $A_{\alpha}^{i}-A_{\alpha}^{*i}$ and $A_{\alpha}^{f}-A_{\alpha}^{f*}$ in the optimal protocol. This implies that the computation of the initial geodesic for the finite-time process is based on the assumption of the absence of kinks at the initial time $t=0$. However, the finite-time optimal protocol $A_{\alpha}^{i*}\to A_{\alpha}^{f*}$ is obtained by shifting this protocol (counter-diabatic driving) to obtain the kink at both $t=0^{+}$ and $t=\tau^{-}$.
This contradicts its formulation of constrained and unconstrained optimization for the initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, respectively. Hence, mathematically, this is interpreted as a boundary condition artifact because of the inability of assigning a thermodynamic cost to kinks within their framework. In contrast, within MinAP formalism, a thermodynamic cost is inherently assigned for kinks, and hence unconstrained optimization is implemented for both initial ($t=0^{+}$) and final ($t=\tau^{-}$) points, signified by a double optimization problem in eq.˜17. Therefore, our optimization framework uses minimum assumptions, equivalently assuming the least information about optimal protocols. This allows us to implement an unconstrained optimization problem in eq.˜17 that searches/accesses a wider control parameter space for the optimal protocol.
These differences lead to rather drastic physical results. For example, the short-time optimal protocol eq.˜23 is midpoint interpolation in the geodesic space $\mathcal{G}(A_{\alpha})$ and not necessarily in control parameter space $A_{\alpha}$ unless $\mathcal{G}(A_{\alpha})=A_{\alpha}$. Hence, even for short driving time $\tau$, the finite-time optimal protocol eq.˜23 respects the Riemannian geometric structure in control parameter space. In contrast, the short-time protocols obtained in Ref. [10, 11] follow a midpoint interpolation in control parameter space. This difference is irrelevant in the case of the optimal control of the trap centre, since $\mathcal{G}(A_{\alpha})=A_{\alpha}$ for the control of the trap centre and the short-time optimal protocols obtained in Ref. [81, 82]. However, $\mathcal{G}(A_{\alpha})\neq A_{\alpha}$ for the optimal control of the trap stiffness [24], giving divergent analytical results compared to Ref. [10, 11, 81, 82]. Subsequently, we will revisit this issue in section˜5.2 with the example of a stochastic particle in a harmonic trap and reconcile the discrepancies between optimal transport theory [49, 50, 58] and thermodynamic geometry [24].
4 Generalizations
4.1 Multiple control parameters
The driving EPR in eq.˜5 assumes that all driven control parameters are decoupled and independently driven. However, in practice, there could be cross-couplings between different control parameters. In addition, the transition affinity only accounts for the anti-symmetric part in the transition space [5], and the transition mobility $D_{\alpha}$ were fixed.
For a slow driving of multiple parameters, the driving Lagrangian is given by:
$\begin{split}\mathcal{L}_{drv}^{*}\left[\lambda\right]=\frac{1}{2}g_{ij}\dot{\lambda}^{i}\dot{\lambda}^{j},\end{split}$ (28)
where eq.˜28 is represented in Einstein’s summation convention, which signifies the summation over the contracted index. $\vec{\lambda}$ is the vector of control parameters with $i^{th}$ component $\lambda^{i}$. The Hessian of $\mathcal{L}^{*}$ (mass) gives the metric tensor $g_{ij}$,
[…]
The MinAP implies solving a variational optimization problem, $\Sigma_{qs}^{*}=\inf(\Sigma_{drv}^{qs})=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$. The geodesic equation for the control parameter that minimizes $\Sigma_{drv}^{qs}$ eq.˜28 reads,
$\begin{split}\ddot{\lambda}^{i}+\Gamma_{jk}^{i}\dot{\lambda}^{j}\dot{\lambda}^{k}=0,\end{split}$ (30)
where the Christoffel symbols are defined as
$\begin{split}\Gamma_{jk}^{i}=\frac{1}{2}g^{im}\left(\frac{\partial g_{mj}}{\partial{\lambda}^{k}}+\frac{\partial g_{mk}}{\partial\lambda^{j}}-\frac{\partial g_{jk}}{\partial\lambda^{m}}\right)\end{split}$ (31)
Equations˜28 and 30 are equivalent to eqs.˜5 and 7 for the multi-parameter slow-driving optimal control problem.
By construction, the geodesic $\mathcal{G}(\vec{\lambda})$ is the solution that minimizes $\Sigma_{drv}^{qs}$, $\mathcal{G}(\vec{\lambda})=\arg\inf_{\{\lambda\}}\left(\Sigma_{drv}^{qs}\right)$. The corresponding optimal protocol is the linear interpolation, which reads,
$\begin{split}\mathcal{G}(\vec{\lambda})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\vec{\lambda}^{inl})+\frac{t}{\tau}\mathcal{G}(\vec{\lambda}^{fnl}).\end{split}$ (32)
To avoid confusion with the vector index, we denote the vectors of initial and final control parameters by $\vec{\lambda}^{inl}$ and $\vec{\lambda}^{fnl}$, respectively. Using eq.˜32, the optimal slow driving EP reads:
[…]
Here, $||*||_{2}$ is the $L^{2}-norm$ that quantifies the distance in the Riemannian manifold induced by multiple control parameters. The finite-time optimal control framework developed in section˜3 is based on the MinAP combined with the existence of the slow-driven geodesic $\mathcal{G}(\vec{\lambda})$ and incorporating the possibility of kinks. Hence, a multi-parameter finite-time optimal control protocol is rather trivial, provided that a slow driving geodesic eq.˜32 exists and is computed. The corresponding finite-time multi-parameter optimal protocol reads,
$\begin{split}\mathcal{G}_{\tau}(\vec{\lambda})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}^{inl})+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}^{fnl}),\end{split}$ (34)
with the finite-time multi-parameter optimal driving EP reads:
[…]
Equations˜32, 33, 34 and 35 are a generalization of eqs.˜8, 9, 21 and 25, respectively, for the multi-parameter generalized optimal control problem.
4.2 Numerical computation of geodesic
We have focused on the exact analytical formulation of the GFTOC framework. However, analytically computing the exact geodesic $\mathcal{G}(\vec{\lambda})$ is not always feasible for more sophisticated problems. To this end, a numerical framework is required to compute the geodesic/optimal protocols. Here, we highlight two different numerical algorithms that are implemented to calculate the geodesic.
The first algorithm is based on the MinAP, which focuses on the thermodynamic cost of driving, where the geodesic $\mathcal{G}(\vec{\lambda})$ is obtained by the minimization of the Lagrangian eq.˜28, and is extended to return the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$, it reads,
Algorithm 1 Computing finite-time optimal protocols by minimizing the cost function (action)
Input: $\vec{\lambda}^{inl},\vec{\lambda}^{fnl},\tau,g_{ij}$, eqs.˜28, 34 and 35
1:Compute the geodesic $\mathcal{G}(\vec{\lambda})$ obtained by the recursive minimization of eq.˜28, and that connects $\mathcal{G}(\vec{\lambda}^{inl})$ and $\mathcal{G}(\vec{\lambda}^{fnl})$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$.
Note that, Algorithm 1 is modified to directly compute $\mathcal{G}_{\tau}(\vec{\lambda})$ instead of $\mathcal{G}(\vec{\lambda})$, if the thermodynamic cost for kinks is assigned under the action minimization problem and the endpoint discontinuity in the protocol space is allowed by relaxing the boundary condition constraint. Algorithm 1 extends to the class of numerical optimization methods developed in Ref.[39, 40, 59, 60, 61, 31, 62, 63, 64, 65, 66] and related to machine learning techniques Ref.[67, 68, 69]. However, none of the previous numerical algorithms incorporates kinks as a result of their failure to assign a thermodynamic cost to kinks. To our knowledge, Algorithm 1 is the novel formulation that numerically computes finite-time optimal protocols with kinks and computes the associated finite-time driving EP cost.
A second alternative algorithm computes the finite-time optimal protocols by numerically solving the geodesic eq.˜30 as a boundary value problem, focusing on driving dynamics. The algorithm for it reads:
Algorithm 2 Computing finite-time optimal protocols by numerically solving geodesic ODEs
Input: $\vec{\lambda}^{inl},\vec{\lambda}^{fnl},\tau,g_{ij},\Gamma_{ij}^{k}$, eq.˜30 and eq.˜34
1:Compute the geodesic $\mathcal{G}(\vec{\lambda})$ by numerically solving eq.˜30 under the boundary value constraint $\vec{\lambda}^{inl}$ and $\vec{\lambda}^{fnl}$ at $t=0$ and $t=\tau$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
3:Return the optimal protocol $\mathcal{G}_{\tau}(\vec{\lambda})$ and EP cost $\Sigma_{\tau}^{*}$.
In contrast to Algorithm 1, Algorithm 2 can only compute $\mathcal{G}(\vec{\lambda})$, and requires eq.˜34 to compute $\mathcal{G}_{\tau}(\vec{\lambda})$. The numerical approach elaborated in this section broadens the applicability of the GFTOC when an exact analytical solution of the geodesic is not available.
5 Applications of GFTOC framework
5.1 Optimal control of free energy
The EP defined for the driving of the free energy $\psi_{E}$ by changing the control parameters $\{\lambda_{E}\}$ of $E$ reads, $\mathcal{L}_{drv}^{*}=\frac{1}{2}\left(\partial^{2}\psi_{E}/\partial{\lambda_{E}^{i}}\partial{\lambda_{E}^{j}}\right)\dot{\lambda}_{E}^{i}\dot{\lambda}_{E}^{j}$. The initial and final control parameter vectors are denoted by $\vec{\lambda}_{E}^{inl}$ and $\vec{\lambda}_{E}^{fnl}$, respectively. The corresponding optimal slow driving EP and the optimal driving protocol read:
$\begin{split}&\Sigma_{qs}^{*}=\frac{1}{2\tau}\left(\mathcal{G}(\vec{\lambda}_{E}^{inl})-\mathcal{G}(\vec{\lambda}_{E}^{fnl})\right)^{2},\\ &\mathcal{G}(\vec{\lambda}_{E})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(\vec{\lambda}_{E}^{inl})+\frac{t}{\tau}\mathcal{G}(\vec{\lambda}_{E}^{fnl}).\end{split}$ (36)
The corresponding optimal finite-time EP and the optimal finite-time driving protocol read:
$\begin{split}\Sigma_{\tau}^{*}&=\frac{1}{2(2+\tau)}\left(\mathcal{G}(\vec{\lambda}_{E}^{inl})-\mathcal{G}(\vec{\lambda}_{E}^{fnl})\right)^{2},\\ \mathcal{G}_{\tau}(\vec{\lambda}_{E})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}_{E}^{inl})+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}(\vec{\lambda}_{E}^{fnl}).\end{split}$ (37)
Note that the exact geodesic functions $\mathcal{G}(\vec{\lambda}_{E})$ are model-specific, due to the different analytical dependence of $\psi_{E}$ on $\lambda_{E}$ [27, 28, 26]. But in the next section, we explicitly compute the geodesic functions $\mathcal{G}(\vec{\lambda}_{E})$ for a stochastic particle in a harmonic trap.
5.2 Stochastic particle in a harmonic trap and
Wasserstein distance for the Gaussian distribution
A stochastic particle in a harmonic trap potential models an important equivalence class that describes a large class of physical systems, and hence has been the testbed of novel theoretical developments in ST. The stochastic particle satisfies the Boltzmann distribution, $P\propto e^{-k(x-m)^{2}}$, with the trap stiffness $k$ and center $m$. In this case, these two relevant control parameters fully capture the statistical properties of the system. We consider a finite-time optimal control problem for the time-dependent change of the trap stiffness $k(t)$ and center $m(t)$ [10, 24]. The inverse of the trap stiffness quantifies the covariance $(C)$ of the Gaussian distribution for the particle position, $C\propto 1/k$ [10]. Hence, by definition, the finite-time optimal control formulation for changing the center and stiffness of the harmonic trap is equivalent to the optimal transport theory to change the mean and covariance of the Gaussian distribution for the position of the particles [49] or the application of thermodynamic geometry to the harmonic trap [24]. However, the results obtained from different methods do not match [10, 24, 49, 58]. We resolve these discrepancies using the GFTOC framework.
We consider an optimal control problem for changing the trap stiffness and center from $\{m_{i},k_{i}\}$ to $\{m_{f},k_{f}\}$. The mass for the center and the stiffness of the trap are $g_{mm}=1$ and $g_{kk}=1/k^{3}$ [24]. This implies the geodesic function $\mathcal{G}(m)=m$ and $\mathcal{G}(k)=k^{-1/2}$. Using $C\propto 1/k$, we recover the geodesic function $\mathcal{G}(C)=C^{1/2}$ for the covariance. Hence, the minimum slow driving EP and the corresponding optimal protocols for the mean and variance read:
$\Sigma_{qs}^{*}=\frac{1}{2\tau}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right],$ (38a)
$m(t)=\left(1-\frac{t}{\tau}\right)m_{i}+\frac{t}{\tau}m_{f},$ (38b)
$\sqrt{C(t)}=\left(1-\frac{t}{\tau}\right)\sqrt{C_{i}}+\frac{t}{\tau}\sqrt{C_{f}}.$ (38c)
Equation˜38a is the square of the $\mathcal{W}_{2}$ Wasserstein distance between the initial and final Gaussian probability distributions divided by $2\tau$ in Refs. [49, 50, 54] with the corresponding optimal protocols for the mean and covariance eqs.˜38b and 38c mentioned in Ref.[49].
Further, applying the GFTOC framework, the minimum finite-time driving EP and finite-time optimal protocols for the mean and variance read:
$\Sigma_{\tau}^{*}=\frac{1}{2(2+\tau)}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right],$ (39a)
$m_{\tau}(t)=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)m_{i}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)m_{f},$ (39b)
$\sqrt{C_{\tau}(t)}=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\sqrt{C_{i}}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\sqrt{C_{f}}.$ (39c)
Equation˜39a is the novel formulation for finite-time processes, which requires dividing the square of $\mathcal{W}_{2}$ Wasserstein distance between the initial and final Gaussian probability distributions by $2(2+\tau)$; which is less than eq.˜38a due to kinks. The finite-time optimal protocol for the center of the trap eq.˜39b is the same as Ref.[10, 11]. However, the finite-time optimal protocol for the trap stiffness eq.˜39c is different from Refs. [10, 11]. These differences, due to boundary conditions and their physical implications, have previously been detailed in section˜3.5.
5.3 Optimal control of excess EP
The optimal control formulation has been developed for the affinity, which lies in the transition space. However, as previously shown in eq.˜2, a part of the EPR can be integrated to obtain the excess EP $\Sigma^{ex}=-D_{E}^{KL}$, which is defined in the state space. The integration is implemented: 1) from the transition space to the state space, and 2) in time. This implies that $\Sigma^{ex}$ is the boundary term that does not require knowledge of the transition-space topology. This property of the excess EP has been exploited in the literature so far to formulate the control or geodesic description of cEQ systems [83, 23, 84, 54, 85, 86, 87]. For $\Sigma^{ex}=\sum_{\{i\}}\rho_{i}\log{(\rho_{i}/\rho_{i}^{E})}$, the mass is equal to the stochastic Fisher information $\partial^{2}_{\rho_{i}}\Sigma^{ex}=1/\rho_{i}$ defined with respect to $\rho_{i}$  [^6] Note that we use the negative of the excess EP to ensure the positivity of the mass.. Reorganization of ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}(\partial_{t}\rho_{i})^{2}/2\rho_{i}$ leads to ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}\rho_{i}\left[\partial_{t}\log{(\rho_{i}/\rho_{i}^{E})}\right]^{2}$, the quadratic form of the excess driving EP written in terms of the driving of the excess affinity $A_{i}^{ex}=-\log(\rho_{i}/\rho_{i}^{E})$. Thus, the control of the state is equivalent to the control of the excess affinity. To avoid confusion with the state index, we introduce a short-hand notation $\vec{\rho}(\tau)$ and $\vec{\rho}(0)$ for the final and initial states of the system represented as component-wise vectors.
Using $\mathcal{L}_{drv}^{ex}=\sum_{\{i\}}(\partial_{t}\rho_{i})^{2}/2\rho_{i}$ leads to $\mathcal{G}(\rho_{i})=2\sqrt{\rho_{i}}$ and resolves $\Sigma_{drv}^{ex}=\int_{0}^{\tau}\mathcal{L}_{drv}^{ex}dt$. The corresponding slow-driving optimal excess EP and the optimal driving protocol read:
$\begin{split}\Sigma_{qs}^{*ex}=\frac{2}{\tau}\left(\sqrt{\vec{\rho}(\tau)}-\sqrt{\vec{\rho}(0)}\right)^{2},\end{split}$ (40a)
$\begin{split}\sqrt{\vec{\rho}(t)}&=\left(1-\frac{t}{\tau}\right)\sqrt{\vec{\rho}(0)}+\frac{t}{\tau}\sqrt{\vec{\rho}(\tau)}.\end{split}$ (40b)
Equation˜40a is the cEQ slow-driving formulation with a quadratic relation between the thermodynamic length and the excess EP [84, 83]. However, the corresponding optimal slow-driving protocol eq.˜40b is novel.
Furthermore, the solution of the novel GFTOC problem for excess EP leads to the optimal finite-time excess EP and the finite-time optimal driving protocol, which read:
$\begin{split}\Sigma_{\tau}^{*ex}&=\frac{2}{(2+\tau)}\left(\sqrt{\vec{\rho}(\tau)}-\sqrt{\vec{\rho}(0)}\right)^{2},\end{split}$ (41a)
$\begin{split}\sqrt{\vec{\rho}_{\tau}(t)}&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\sqrt{\vec{\rho}(0)}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\sqrt{\vec{\rho}(\tau)}.\end{split}$ (41b)
Equation˜41a reveals that the ‘thermodynamic shock’ lowers the driving cost of excess EP due to kinks in $\sqrt{\vec{\rho}_{\tau}(t)}$. Within infinitesimal time at the initial and final driving times, the state of the system undergoes an instantaneous simultaneous jump. Importantly, eqs.˜40 and 41 are constrained by the conservation law manifold. Therefore, they can be further simplified/sophisticated to obtain eqs.˜40 and 41 as a function of the linearly independent state space whose dimension is reduced exactly by the total number of conservation laws.
Since, the choice of reference gauge $E$ itself is flexible and should be defined according to physical/experimental constraints, this formulation extends to any reference gauge, for example, $E=ss$ corresponds to any non-equilibrium steady state with a focus on the optimization of the excess EP.
5.4 Linear optimal control of housekeeping EPR
Linear irreversible thermodynamics utilizes the most fundamental quadratic dissipation function $\mathcal{L}^{*}=kA_{\alpha}^{2}$ for the EPR [89, 90], which is also the regime associated with the linear response theories in non-equilibrium thermodynamics. This leads to $\mathcal{L}_{drv}^{*}=k\dot{A}_{\alpha}^{2}$ with the mass $\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}=2k$. Fixing $k=1/2$ implies a unit mass. Solving $\dot{A}_{\alpha}=v_{qs}$ gives $\mathcal{G}^{lin}(A_{\alpha})=A_{\alpha}$. Thus, the geodesic is a linear interpolation between $A_{\alpha}^{i}$ and $A_{\alpha}^{j}$ in $A_{\alpha}$ space, and the slow driving optimal EP and the corresponding optimal protocol read,
$\begin{split}\Sigma_{qs}^{*}&=\frac{1}{2\tau}\left(A_{f}-A_{i}\right)^{2},\\ A_{\alpha}(t)&=\left(1-\frac{t}{\tau}\right)A_{\alpha}^{i}+\frac{t}{\tau}A_{\alpha}^{f},\end{split}$ (42)
respectively. The optimal finite-time driving EP and the corresponding optimal finite-time protocol read,
$\begin{split}\Sigma_{\tau}^{*}&=\frac{1}{2(2+\tau)}\left(A_{f}-A_{i}\right)^{2},\\ A_{\alpha}(t)&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)A_{\alpha}^{i}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)A_{\alpha}^{f}.\end{split}$ (43)
The linear optimal control formulated here is a trivial example of Euclidean geometry with a constant mass. This also highlights how, for the cEQ regimes where housekeeping EPR becomes substantially dominant over free energy and excess EP, a trivial linear interpolation between the control parameters leads to a better optimal control procedure than in sections˜5.1 and 5.3.
5.5 Optimal control of housekeeping EPR
Using the housekeeping EPR in eq.˜2, we formulate a novel and up-to-unresolved optimal control problem for $\dot{\Sigma}_{hk}$, using the external control of $F_{\alpha}$ from an initial to a final value $F_{\gamma}^{i}$ to $F_{\gamma}^{f}$. Physically, this implies that the control parameters $\{\lambda_{E}\}$ of $E$ are fixed and the state-space distribution is in steady state. This leads to vanishing driving work and excess EPR contributions, respectively. The driving Lagrangian for the EPR of housekeeping then reads $\mathcal{L}_{drv}^{*}=\frac{1}{2}\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}\left(\dot{F}_{{\alpha}}\right)^{2}$. The mass for the driving reads, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}^{\perp}\cosh{\left(F_{\alpha}/2\right)}+T_{\alpha}^{\perp}F_{\alpha}\sinh{\left(F_{\alpha}/2\right)}/4$. It is rewritten as $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=\frac{1}{2}\left(T_{\alpha}(A_{\alpha})+T_{\alpha}(A_{\alpha}^{\dagger})\right)+\frac{1}{4}\dot{\Sigma}_{\alpha}^{hk}$ with the total affinity $A_{\alpha}=F_{\alpha}+A_{\alpha}^{ex}$ and $A_{\alpha}^{\dagger}=F_{\alpha}-A_{\alpha}^{ex}$ being the affinity obtained by the time-reversal of the boundary term. Using the orthogonal symmetry of the non-equilibrium fluctuations, $T_{\alpha}(A_{\alpha})=T_{\alpha}(A_{\alpha}^{\dagger})$, the mass is reduced to, $\partial_{{F_{\alpha}}}^{2}\mathcal{L}_{hk}^{*}=T_{\alpha}+\dot{\Sigma}_{\alpha}^{hk}/4$. Hence, for the fEQ driving of $F_{\alpha}$, the mass is proportional to the traffic $T_{\alpha}$ and the housekeeping EPR $\dot{\Sigma}_{\alpha}^{hk}$ due to the bidirectional transition $\alpha$, and increases exponentially as the system is further from equilibrium due to the hyperbolic scalings of $T_{\alpha}$ and $\dot{\Sigma}_{\alpha}$. This reveals that the higher housekeeping EPR and non-equilibrium fluctuations generate higher driving resistance. Therefore, we use an exponential approximation of the hyperbolic functions, which is valid for fEQ systems, and use a previously computed closed-form analytical expression for the fEQ geodesic $\mathcal{G}^{fEQ}$ in eq.˜10.
The solution of the slow-driving optimal control problem for the housekeeping EPR leads to the optimal driving EP and the corresponding optimal protocol,
$\begin{split}\Sigma_{qs}^{*hk}&=\frac{1}{2\tau}\left(\mathcal{G}^{fEQ}(F_{\alpha}^{f})-\mathcal{G}^{fEQ}(F_{\alpha}^{i})\right)^{2},\\ \mathcal{G}^{fEQ}(F_{\alpha})&=\left(1-\frac{t}{\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{i})+\frac{t}{\tau}\mathcal{G}^{fEQ}(F_{\alpha}^{f}).\end{split}$ (44)
Similarly, the solution of the finite-time optimal control problem for the housekeeping EPR leads to the optimal finite-time driving EP and the corresponding finite-time optimal protocol, which are,
$\begin{split}\Sigma_{\tau}^{*hk}&=\frac{1}{2(2+\tau)}\left(\mathcal{G}^{fEQ}(F_{\alpha}^{f})-\mathcal{G}^{fEQ}(F_{\alpha}^{i})\right)^{2},\\ \mathcal{G}_{\tau}^{fEQ}(F_{\alpha})&=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{i})\\ &\hskip 28.45274pt+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\mathcal{G}^{fEQ}(F_{\alpha}^{f}).\end{split}$ (45)
A pictorial representation of the optimal control of $F_{\alpha}$ is shown in fig.˜2 and pictorially summarizes the essence of this work.
5.6 Multi-parameter optimal control of
affinity and mobility
We show the applicability of the multi-parameter optimal control. We incorporate the symmetric part of the transition, namely mobility $D_{\alpha}$, which was previously fixed. This implies $\{{\lambda}_{\alpha}^{i}\}=\{A_{\alpha},D_{\alpha}\},\forall\alpha\in\{\gamma^{\rightleftharpoons}\}$. Henceforth, we use the shorthand notation $a=A_{\alpha}$ and $d=D_{\alpha}$ for brevity. Using $\mathcal{L}^{*}$, we pose a novel finite-time optimal control problem for the simultaneous control of mobility $D_{\alpha}$ and affinity $A_{\alpha}$, Then, the exact metric tensor reads
[…]
Using eqs.˜46 and 31, the exact expression for the Christoffel symbols reads,
$\begin{split}\Gamma_{aa}^{a}&=\left(1+\frac{a}{4}\tanh{\left(\frac{a}{2}\right)}\right)/\left(a+2\tanh{\left(\frac{a}{2}\right)}\right),\\ \Gamma_{da}^{d}&=\Gamma_{ad}^{d}=\left(1+\frac{a}{4}\tanh{\left(\frac{a}{2}\right)}\right)/\left(a+2\tanh{\left(\frac{a}{2}\right)}\right),\\ \Gamma_{aa}^{d}&=d(-14+a^{2}-2\cosh{(a)})/\left(8\left[a\cosh{\left(\frac{a}{2}\right)}+2\sinh{\left(\frac{a}{2}\right)}\right]^{2}\right),\\ \Gamma_{ad}^{a}&=\Gamma_{da}^{a}=\Gamma_{dd}^{a}=\Gamma_{dd}^{d}=0.\end{split}$ (47)
The exact geodesic eq.˜30 for the optimal control of transition affinity and mobility is reduced to,
$\ddot{A}_{\alpha}+\Gamma_{aa}^{a}\>\dot{A}_{\alpha}^{2}=0,$ (48a)
$\ddot{D}_{\alpha}+\Gamma_{aa}^{d}\>\dot{A}_{\alpha}^{2}+2\Gamma_{ad}^{d}\>\dot{A}_{\alpha}\dot{D}_{\alpha}=0.$ (48b)
Equation˜48a for the optimal driving of the affinity is the same as eq.˜7, and does not depend on $D_{\alpha}$. Therefore, the optimal driving dynamics of the mobility $D_{\alpha}$ given by eq.˜48b are enslaved by the geodesic for $A_{\alpha}$. This implies that an optimal control problem for transition affinity and mobility can be reduced to two steps. First step, to solve the optimal control problem $A_{\alpha}$, as formulated in sections˜2 and 3, and obtain the geodesic $\mathcal{G}(A_{\alpha})$. Second step, plugging $\mathcal{G}(A_{\alpha})$ into eq.˜48b and solving ODE eq.˜48b for optimal driving of $D_{\alpha}$, and obtain the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$. We cannot find an exact analytical solution for the geodesic $\mathcal{G}(A_{\alpha},D_{\alpha})$ due to the sophistication associated with solving coupled ODEs. However, the numerical computation of $\mathcal{G}(A_{\alpha},D_{\alpha})$ is always feasible.
6 Conclusion and Outlook
We consider a finite-time optimal control problem for driving the control parameters of discrete-state systems from the initial to the final value, such that the EPR for driving is minimised. Building upon the ‘minimum action principle’ (MinAP), we propose a framework for the generalized finite-time optimal control (GFTOC). The ‘minimum action principle’ allows for a variational formulation of the finite-time optimal control problem. To solve this problem, we develop a two-step solution. In the first step, assuming slow driving, we solve the slow driving optimal control problem by exploiting the Riemannian geometry induced in the manifold of control parameters. This approach unifies and generalizes the thermodynamic geometry and the Optimal TT for fEQ systems. Then, the dissipation-minimizing optimal paths are obtained by the geodesic, the minimum-distance path on the Riemannian manifold.
In the second step, we investigate the possibility of discontinuous endpoint jumps in the optimal protocol by relaxing the assumption of slow driving. We utilize the slow-driving geodesic obtained to compute the finite-time geodesic; our analysis reveals an exact mapping between the slow-driving and finite-time optimal control problem. This strong and important result allows us to extend the framework of thermodynamic geometry developed for slowly driven systems to systems driven in any finite time. Importantly, due to the ‘minimum action principle’, a thermodynamic cost is assigned to the endpoint jumps; this instantaneous dissipation cost at the endpoints is termed a ‘thermodynamic shock’. We analytically prove that it is a model-independent generic mechanism for far-from-equilibrium (fEQ) systems, which reduces the driving dissipation cost. Within our framework, discontinuous endpoint jumps in the optimal protocol are a physical manifestation of a ‘thermodynamic shock’, implying that the optimal driven dynamics in finite-time are constrained by thermodynamics.
Our framework opens up a plethora of practical applications in biology, chemistry, population dynamics, and nanoscale / mesoscale devices, where stochastic thermodynamics has been an experimentally tested theoretical paradigm [4]. Our framework opens the door for the experimental design, optimization, and control of such systems. The experimental verification of our framework awaits exploration [37]. The GFTOC framework is rather trivially extended to quantum systems, as the development of optimal control and speed limits for quantum systems is attributed to the underlying geometric structure and not to the quantum nature [91, 92, 93]. More broadly, the GFTOC methodology developed here in stochastic thermodynamics applies to other interdisciplinary domains that rely on Riemannian geometry as their mathematical foundation, including information geometry [38], in which the role of ‘kinks’ warrants further investigation.
Acknowledgements.
ATM thanks Jin-Fu Chen for pointing out the experimental work Ref.[37], which motivated the theoretical formulation of the GFTOC problem.
Note added. —  After the first publicly available draft of this work was released, we noted that the GFTOC framework solves several major open problems identified in the ‘Perspectives and Conclusions’ section of the recent review Ref.[12].